OpenStudy (anonymous):

Simplify ³√-243a^6b^12

OpenStudy (anonymous):

go on mathway.com and it will simplify that for you quickly

OpenStudy (tkhunny):

Or, if you wanted to learn something, you could wade through it on your own! \(\sqrt[3]{-243\cdot a^{6}\cdot b^{12}} = (-243)^{1/3}\cdot a^{6/3}\cdot b^{12/3}\) Now what?